PROVE THAT - tan A+B/2=cot C/2

PROVE THAT -     tan A+B/2=cot C/2

Grade:

1 Answers

ruchi yadav
10 years ago
A +B +C =180

tan(A+B/2) = tan (180-C/2) = tan(90-<C/2>)
which is equal to cot C/2

Thank You
Ruchi
Askiitians Faculty